# Transfer learning with generative models for object detection on limited datasets

### v1

@@ 7 Conclusion
We have proposed a method that leverages a generative model pretrained on extensive and diverse datasets, to generate images for the pretraining of an object detector and subsequent transfer learning to the real domain. We have demonstrated remarkable results on both NuImages and OzFish datasets. We have showed that our method allows to achieve good detection performance with just few hundreds of real images, comparable to training the model on thousands of real images. Furthermore, our method significantly reduces the human effort involved in the annotation process, limiting the need for labor-intensive labeling tasks.
During the preparation of our manuscript, an independent paper has appeared online as preprint in Ref. [7], exploring the potential of training an object detector using generated images. However, their generative model was trained on a substantial amount of real data from a specific dataset, thus without addressing the issue of limited datasets from a generic domain that we are considering here. Besides, while via their model (GeoDiffusion) they primarily report the FID as the standard metric to evaluate generative models, we here also explore how FID impacts on mAP. While our findings are in their preliminary stages and comprise hypotheses, they open new possibilities for understanding the relationship between FID and mAP, hence providing an intriguing direction for further exploration and investigation.
@@
In conclusion, our method offers a novel and efficient solution to tackle the data scarcity challenge in object detection, paving the way for further exploration of generative models applications in this field. This can be useful in specific domains where the gathering and labeling of data can be problematic or costly, such as, among others, medicine and molecular biology, atmospheric monitoring, material science, space imaging, and non-human-friendly environments.

### v2

@@ 7 Conclusion
We have proposed a method that leverages GLIGEN, a generative model pretrained on extensive and diverse datasets, to generate images with targets inside bounding boxes for the pretraining of an object detector and subsequent transfer learning to the real domain. We have demonstrated remarkable results on both NuImages and OzFish datasets. We have showed that our method allows to achieve good detection performance with just few hundreds of real images, comparable to training the model on thousands of real images. Furthermore, our method significantly reduces the human effort involved in the annotation process, limiting the need for labor-intensive labeling tasks.
In our work we address the generalization capability of the proposed approach by comparing a common-domain dataset of cars in urban environment with a more specific and uncommon dataset of fishes in underwater environment. Besides, it is plausible that our framework can be generalizable even to more complex scenarios. Indeed, GLIGEN leverages CLIP, i.e. a model trained on vast and diverse datasets that encompass categories beyond just underwater fish and urban cars. CLIP requires only image-text pairs, which are more easy to obtain compared to labor-intensive annotations like bounding boxes. GLIGEN builds upon CLIP’s strong representation capabilities and learns only the additional ability to depict objects within the desired bounding boxes. This separation of capabilities suggests that the two aspects are relatively independent. Therefore, in more niche scenarios, such as on medical images, it should be sufficient to fine-tune the underlying CLIP model [69].
During the preparation of our manuscript, an independent paper has appeared online as preprint in Ref. [45], exploring the potential of training an object detector using generated images. However, their generative model was trained on a substantial amount of real data from a specific dataset, thus without addressing the issue of limited datasets from a generic domain that we have considered here. Besides, while via their model they primarily report the FID as the standard metric to evaluate generative models, we here also explore how FID impacts on mAP. While our findings are in their preliminary stages and comprise hypotheses, they open new possibilities for understanding the relationship between FID and mAP, hence providing an intriguing direction for further exploration and investigation.
@@
In conclusion, our method offers a novel and efficient solution to tackle the data scarcity challenge in object detection without the need for manual labeling of data, with the help of a pretrained generative model. In this way, we pave the way for further exploration of generative models applications in this field. This can be useful in specific domains where the gathering and labeling of data can be problematic or costly, such as, among others, medicine and molecular biology, atmospheric monitoring, material science, space imaging, and non-human-friendly environments.
@@ 8 Acknowledgments
Appendix A
The standard ResNet-50 model starts with an initial convolutional layer with 64 $7\times 7$ filters, followed by a $3\times 3$ max pooling layer. The main body of ResNet-50 consists of a series of residual blocks, each functioning as a bottleneck. Each bottleneck comprises a $1\times 1$ convolutional layer to reduce dimensions, a $3\times 3$ convolutional layer to extract spatial features, and another $1\times 1$ convolutional layer to restore the original dimensions, totaling 3 convolutional layers per block. The architecture contains 16 bottleneck blocks arranged in four stages, for a total of 48 convolutional layers: the first stage has 3 blocks with 256 channels, the second stage has 4 blocks with 512 channels, the third stage has 6 blocks with 1024 channels, and the final stage has 3 blocks with 2048 channels. The network concludes with a global average pooling layer followed by a fully connected layer for classification.
In the Faster R-CNN framework, the ResNet50 backbone’s final fully connected layer is omitted. Instead, the outputs of the four intermediate residual blocks, before each upsampling and before the final fully connected layer, are used as inputs for a Feature Pyramid Network (FPN[70]. The FPN is composed of two convolutional layers for each input: a lateral connection with $1\times 1$ filters to reduce the number of channels to 256, and a top-down pathway with a $3\times 3$ convolutional layer with 256 filters. The lateral connections process each input map, while the top-down pathway refines the feature maps, which are combined with the lateral outputs via summation. The final feature map is processed with another $3\times 3$ convolutional layer and a $1\times 1$ layer that proposes the regions of interest.
Two-stage detectors like Faster R-CNN operate in two phases: first generating region proposals (potential bounding boxes) and then refining and classifying these regions. By contrast, FCOS is a single-stage detector directly predicting bounding boxes and class scores from feature maps in a single pass. FCOS eliminates the need for anchor boxes and proposals, streamlining the detection process. It uses a fully convolutional approach to predict bounding box coordinates for each pixel in the feature map. Key components of FCOS include a FPN that generates feature maps at multiple scales with strides of 8, 16, 32, 64, and 128. Each level of the feature pyramid passes through shared heads comprising three parallel branches: classification, bounding box regression, and centerness. The classification branch outputs class scores, while the regression branch predicts the offsets for the bounding boxes. The centerness branch predicts a score indicating how close a pixel is to the center of an object, helping to down-weight low-quality bounding boxes far from object centers.
The MobileNetV3-Large architecture consists of a series of convolutional layers and bottleneck blocks designed for efficient mobile performance. It begins with a standard 3x3 convolutional layer with 16 output channels, followed by 16 bottleneck blocks with varying expansion sizes and kernel sizes, alternating between 3x3 and 5x5 kernels. Some of these blocks incorporate the Squeeze-and-Excite module, which enhances the network’s representational power by adaptively recalibrating channel-wise feature responses through global average pooling and fully connected layers, culminating in a sigmoid activation. Following the bottleneck layers, the architecture includes a 1x1 convolutional layer with 960 output channels, a 7x7 pooling layer, and another 1x1 convolutional layer without batch normalization, expanding to 1280 output channels.